Each of the four engines can produce a power id 56megawatts.assuming no air resistance,how long would it take the plane to reach its cruising speed and altitude?assume the mass of the plane is constant. Give me Answer in minutes and to be one decimal place of accuracy.


1
Expert's answer
2021-06-09T17:54:10-0400
"Pt=mgh+0.5mv^2\\4\cdot56\cdot10^6(60\cdot t)=550\cdot10^6(9.8)(13100)\\+0.5(550\cdot10^6)(305)^2\\t=7000\ min"